Im not really sure how to do this can someone solve these and provide explanation?... 2x+5y=11 6x+4y=11
6x+2Y=-6 8x-y=2

Answer by richwmiller(17219)

2x+5y=11
6x+4y=11
intersect at
x = 1/2, y = 2
6x+2y=-6
8x-y=2
intersect at
x = -1/11, y = -30/11
I am not sure what you have.
Are there two sets of two equations or one set of 4 equations?
None of them are parallel nor perpendicular.
No three of them meet at any one point.